# Lin Li1,2∗,  Wei Chen1∗,  Jiahui Li3,  Long Chen1†

### v1

@@ 4.1 Experimental Setup
4.1.1 Datasets
PSG. The Panoptic Scene Graph (PSG) dataset [4] is built upon the COCO dataset [33], comprising 48,749 annotated images of which 46,563 are for training and 2,186 are for testing. It integrates 80 “thing” object categories and 53 “stuff” object categories, alongside 56 relation categories. To enable open-ended scene graph generation, we followed ASMv2 [1], using scene graph caption where objects are tagged with <ref> and paired with bounding box coordinates specifying their locations, while predicates are marked with <pred> and linked to two bounding boxes indicating their subject and object entities for training.
SWiG. The SWiG dataset [5] extends the imSitu dataset [34] by adding extra bounding box annotations for visible semantic roles, with 69.3% of semantic roles localized. Each image is annotated with a verb and a variable number of semantic roles (1$\sim$6 roles per image), where each verb is associated with three verb frames annotated by independent human annotators. The dataset comprises 25,200 testing images spanning 504 verb categories, 190 semantic role categories, and 9,929 noun entity categories. Unlike traditional output structures that merely predict categories and bounding boxes, we constructed grounded situation frames by filling verb-centric templates (e.g., “The AGENT drinks a LIQUID from a CONTAINER”) with three critical components: role tags (e.g., <agent>, entity categories (e.g., child), and coordinates of corresponding bounding boxes in <box> tags.
4.1.2 Evaluation Metrics
Binary Relation Detection. Following the standard SGG evaluation protocol [15, 1], we reported 1) Recall: The ratio of correctly predicted triplets to the total number of ground-truth triplets, requiring both label accuracy and bounding box IoU >= 0.5 for subject and object. 2) mean Recall (mRecall): The average recall across all predicate categories.
N-ary Relation Detection. We adopted the same five evaluation metrics as [5]: 1) verb: the accuracy of verb prediction. 2) value: the noun accuracy for individual semantic role. 3) value-all (val-all): Overall noun prediction accuracy across all semantic roles in an event. 4) grounded-value (grnd): Bounding box accuracy for each semantic role, requiring an IoU >= 0.5 between predicted and ground-truth boxes. 5) grounded-value-all (grnd-all): entire bounding box accuracy across all semantic roles in an event.
4.1.3 Implementation Details
We utilized Qwen2.5-VL-3B [31] as our base model. Relation-R1 was trained on an 8xA100 GPU server. During the SFT phase, we employed a total batch size of 64, with 8 samples processed per training step. For the GRPO component, we used a total batch size of 32. The learning rates were set to 1e-5 for SFT and 1e-7 for GRPO, respectively.
Table 1: Performance(%) comparison with various SGG models on the PSG dataset [4].
            Method Model Size Answer Recall mRecall Mean
IMP [35] ${}_{\text{CVPR'17}}$ Close-ended 16.50 6.50 11.50
@@ 4.1.3 Implementation Details
TextPSG [38] ${}_{\text{ICCV'23}}$ Open-ended 4.80 - -
ASMv2 [1] ${}_{\text{ECCV'24}}$ 13B 14.20 10.30 12.23
@@
         Method Answer Verb Value Value-all Gnd Value Gnd Value-all
@@
Relation-R1 (Ours) 53.76 43.55 28.23 36.79 26.84
Table 3: Few-Shot GSR performance comparison. $\dagger$ denotes employing the SFT phrase without CoT.
         Method Setting Verb Value Value-all Gnd Value Gnd Value-all
@@
(a) SGG.
(b) GSR.
Figure 3: The statistics of rewards during training.
Figure 4: Performance(%) comparison between ASMv2 and Relation-R1 over all predicates on test set of PSG. The orange area denotes the predicate distribution of the training set.
Figure 5: Qualitative results of binary relation detection (scene graph generation).
Figure 6: Qualitative results of N-ary relation detection (grounded situation recognition).
4.2 Quantitative Comparison Result
4.2.1 Binary Relation Detection
Open-ended SGG. We evaluated Relation-R1 on PSG, comparing it against various established and recent SGG models, as detailed in Table 1. Positioned within the open-ended generation category, Relation-R1 achieved SOTA performance with the highest mRecall of 18.30% and overall Mean of 18.95%, substantially surpassing prior open-ended methods like LLaVA-SpaceSGG (13.23% mRecall, 14.33% Mean). It also achieved a competitive Recall of 19.59%. Notably, Relation-R1 attained these leading results using a significantly smaller 3B parameter model compared to the 13B models used by recent competitors like ASMv2 and LLaVA-SpaceSGG. In addition, we visualized predicate-level comparison against ASMv2 and the predicate distribution in Figure 4. As seen, Relation-R1 outperforms ASMv2 by a clear margin on both prevalent and rare predicates. These results clearly validated the effectiveness and efficiency of our approach for the SGG task in the open-ended setting.
4.2.2 N-ary Relation Detection
Open-ended GSR. We evaluated Relation-R1 for N-ary relation detection on SWiG, comparing it against prior work in Table 2. Within the open-ended setting, Relation-R1 established a new SOTA, achieving top performance across all reported metrics: Verb (53.76%), Value (43.55%), Value-all (28.23%), Gnd Value (36.79%), and Gnd Value-all (26.84%). Notably, Relation-R1 significantly surpassed the previous best open-ended model, OpenSU [39], exhibiting substantial gains, particularly with a remarkable +11.14% absolute improvement on the challenging Gnd Value-all metric. This clearly demonstrated Relation-R1’s superior capacity for comprehending and grounding complex N-ary relationships in visual scenes.
@@ 4.2.2 N-ary Relation Detection
Few-Shot GSR. We also evaluated Relation-R1’s few-shot generalization capability (Table 3). In the one-shot setting, Relation-R1 achieved dramatic improvements over prior work, reaching 31.67% (Verb) and 3.41% (Gnd Value-all) compared to CoFormer’s [28] 0.63% and 0.01%, respectively. This advantage is maintained in the 4-shot setting, where Relation-R1 reached 38.02% (Verb) and 5.92% (Gnd Value-all), significantly surpassing previous methods. These results clearly demonstrate Relation-R1’s superior ability to generalize and perform robust N-ary relation detection, setting a new SOTA for few-shot GSR.
4.3 Qualitative Result
@@
Figure 5 and Figure 6 provided qualitative illustrations of Relation-R1. For Binary Relation Detection (Figure 5), the visualized thinking process reveals how Relation-R1 effectively decomposes the complex task into constituent cognitive steps: object classification, localization, and relation inference. For instance, after identifying the central bed, Relation-R1 pinpoints related objects like the rug and chair and explicitly grounds the inferred spatial relationships (on, beside) between them. This systematic approach demonstrates the successful decomposition of the scene into accurate, grounded object-relation pairs. Similarly, for N-ary Relation Detection (Figure 6), Relation-R1 showcases its understanding by first identifying the core action (crowning) and then methodically grounding each associated semantic role – Agent (woman), Object (tiara), and Recipient (woman) – to its corresponding entity and bounding box. This structured analysis underscores Relation-R1’s ability to parse complex events involving multiple participants and their distinct roles within the scene.
4.4 Diagnostic Experiment
To gain more insights into Relation-R1, we conducted a set of comprehensive studies on SGG and GSR.
Reward Statistics. We visualized the variation curves of rewards in the SGG task and the GSR task, respectively, as displayed in Figure 3(a) and Figure 3(b). It can be seen that at the beginning stage, both tasks can obtain decent rewards due to the foundation laid by SFT. With the gradual progress of GRPO, the rewards of both tasks show a gradually increasing trend.
@@ 5 Conclusion
This paper reveals critical limitations of existing MLLMs in visual relation understanding, particularly concerning N-ary relationships and the detrimental effects of language-induced bias. We propose Relation-R1, an advanced, unified framework designed to jointly tackle both binary and N-ary relational comprehension. Our core innovation lies in a two-stage approach: first establishing foundational reasoning and structured outputs via cognitive CoT-guided SFT, and subsequently refining the model’s outputs using RL with GRPO. This RL stage, guided by carefully designed multi-rewards, explicitly promotes visual-semantic grounding and robust generalization. Extensive experiments conducted on the challenging PSG and SWiG datasets validate the effectiveness of Relation-R1, demonstrating SOTA performance across both binary and N-ary relation detection tasks. By successfully integrating CoT-guided SFT with R1-inspired RL refinement, Relation-R1 represents a solid step towards equipping MLLMs with more human-like relational reasoning capabilities.

### v2

@@ 4.1 Experimental Setup
Datasets. 1) Panoptic Scene Graph (PSG): The official PSG dataset [17], built upon the COCO dataset [41], comprising 48,749 annotated images of which 46,563 are for training and 2,186 are for testing. It integrates 80 “thing” object categories and 53 “stuff” object categories, alongside 56 relation categories. Notably, ASMv2 [8] reconstructs the dataset into a scene graph caption format, utilizing a split of 42,250 training images and 1,000 evaluation images. To mark fair comparison, we followed ASMv2 [8], using scene graph caption where objects are tagged with <ref> and paired with bounding box coordinates specifying their locations, while predicates are marked with <pred> and linked to two bounding boxes indicating their subject and object entities for training. Besides, we also present results using the conventional scene graph format, i.e., a list of visual triplets structured as [subject, subject bounding box, object, object bounding box, relation] on the official dataset [^1] More details of the scene graph format are in the appendix §C.. 2) SWiG: The SWiG dataset [10] extends the imSitu dataset [42] by adding extra bounding box annotations for semantic roles, with 69.3% of semantic roles localized. Each image is annotated with a verb and a variable number of semantic roles (1$\sim$6 roles per image), where each verb is associated with three verb frames annotated by independent human annotators. The dataset comprises 25,200 testing images spanning 504 verb categories, 190 semantic role categories, and 9,929 noun entity categories. Unlike traditional output structures that merely predict categories and bounding boxes, we constructed grounded situation frames by filling verb-centric templates (e.g., “The AGENT drinks a LIQUID from a CONTAINER”) with three critical components: role tags (e.g., <agent>, entity categories (e.g., child), and coordinates of corresponding bounding boxes in <box> tags.
Evaluation Metrics. For binary relation detection, we followed the standard SGG evaluation protocol [24, 8] 1) Recall: The ratio of correctly predicted triplets to the total number of ground-truth triplets, requiring both label accuracy and bounding box IoU >= 0.5 for subject and object. 2) mean Recall (mRecall): The average recall across all predicate categories. As for N-ary relation detection, we adopted the same five evaluation metrics as [10]: 1) verb: the accuracy of verb prediction. 2) value: the noun accuracy for individual semantic role. 3) value-all (val-all): Overall noun prediction accuracy across all semantic roles in an event. 4) grounded-value (grnd): Bounding box accuracy for each semantic role, requiring an IoU >= 0.5 between predicted and ground-truth boxes. 5) grounded-value-all (grnd-all): entire bounding box accuracy across all semantic roles in an event.
Implementation Details. Due to space constraints, details are provided in the appendix §A.
[…]
         Method Model Size Format Recall mRecall Mean
Close-ended
IMP [16] ${}_{\text{CVPR'17}}$ 16.50 6.50 11.50
@@
Open-ended
TextPSG [45] ${}_{\text{ICCV'23}}$ Standard Scene Graph 4.80 - -
R1-SGG [46] ${}_{\text{ArXiv'25}}$ 2B 27.83 17.03 22.43
R1-SGG [46] ${}_{\text{ArXiv'25}}$ 7B 28.77 17.55 23.16
Relation-R1 (Ours) 3B 25.87 21.32 23.60
ASMv2 [8] ${}_{\text{ECCV'24}}$ 13B Scene Graph Caption 14.20 10.30 12.23
@@
         Method Answer Verb Value Value-all Grnd Grnd-all
@@
Relation-R1 (Ours) 57.26 46.66 30.92 40.21 30.18
4.2 Comparison with State-of-the-Art Methods
We compared our Relation-R1 for unified relation comprehension with the existing state-of-the-art binary relation detection and N-ary relation detection approaches in Table 1 and Table 2, respectively.
[…]
4.3 Diagnostic Experiment
To gain more insights and reduce the mutual influence between the two tasks during training, we conducted a set of comprehensive studies on the Relation-R1 trained with binary and N-ary relation detection, respectively.
Table 3: Performance comparison (%) across various cognitive CoT strategies trained on single binary relation detection and $N$-ary relation detection, separately. Blue indicates metrics without correct verb constraints.
  Binary Relation $N$-ary Relation
Method CoT Recall mRecall Verb Value Value-all Grnd Grnd-all
SFT - 14.83 13.86 56.64 42.65 (60.45) 22.11 (25.21) 37.70 (54.05) 16.35 (19.02)
SFT + RL Template-based 20.24 17.31 58.38 47.75 (66.74) 32.00 (35.67) 41.27 (55.88) 31.16 (38.29)
SFT + RL MLLM-generated 20.66 20.30 53.00 42.27 (63.86) 26.67 (30.79) 35.69 (52.43) 25.89 (33.95)
SFT + RL Progressive 22.57 20.57 71.04 61.26 (78.19) 44.98 (49.73) 49.43 (61.25) 36.09 (42.35)
Cognitive Chain-of-Thoughts. Table 3 compares different cognitive CoT strategies for binary and N-ary relation detection. As seen, introducing RL with various CoT strategies (SFT + RL) demonstrates clear benefits. The template-based CoT improves binary relation Recall to 20.24% and mRecall to 17.31%. For N-ary relations, it enhances performance across the board compared to the constrained SFT baseline, for instance, achieving 47.75% on Value and 31.16% on Grnd-all. The MLLM-generated CoT slightly outperforms the template-based approach in binary relation detection, but underperforms in N-ary relation detection. This is because N-ary relation detection metrics are often limited by verb accuracy, the model may generate a verb different from the GT yet still semantically plausible. Notably, when the verb constraint is removed, MLLM-generated CoT yields significant gains, particularly in Grnd-all (+14.93%). While template-based and MLLM-generated CoTs offer improvements over SFT, the progressive strategy stands out. This approach significantly outperforms all others across all metrics, achieving 22.57% Recall in binary relation detection and top scores (e.g., 71.04% Verb accuracy) in N-ary relation detection. These results highlight the superior ability of the progressive CoT to guide complex relation detection.
Figure 3: The statistics of binary relation rewards.
Figure 4: The statistics of $N$-ary relation rewards.
Reward Analysis. We visualized the reward variation curves for binary and N-ary relation detection in Figure 3 and Figure 4, respectively. It can be seen that at the beginning stage, both tasks can obtain decent rewards due to the foundation laid by SFT. With the gradual progress of GRPO, the rewards of both tasks show a gradually increasing trend. In binary relation detection, guided with progressive CoT yields a slightly enhanced overall reward trajectory, although the improvement is not significant. This might be since the gap between the CoT generated by MLLM (e.g., Qwen [39]) and that produced by the template is not obvious. Conversely, N-ary relation detection exhibits a significant surge in reward, primarily attributed to the model’s emergent capability to generate synonymous grounded situation frames (c.f. §5).
Completion Length Analysis.
Figure 5: The statistics of binary relation completion lengths.
Figure 6: The statistics of $N$-ary relation completion lengths.
The evolution of completion length during GRPO training is depicted in Figure 5 and Figure 6. Consistent with the reward trends, the completion lengths for binary relation detection exhibit a gradual increase as GRPO progresses, suggesting the generation of more detailed and descriptive outputs over time. Notably, the template-based CoT strategy yields significantly shorter completions compared to the more flexible MLLM-generated and progressive CoT strategies, since only the cognitive process is involved. The progressive CoT guidance, by enabling the model to explore synonymous relational expressions and incorporate more nuanced details, leads to an increase in the average completion length.
Table 4: Few-Shot GSR performance comparison.
         Method Setting Verb Value Value-all Grnd Grnd-all
@@ 4.3 Diagnostic Experiment
$w$/$o$ RL 34.88 17.14 3.94 15.63 3.38
@@
Few-Shot N-ary Relation Detection. We also evaluated Relation-R1’s few-shot generalization capability (Table 4). In the one-shot setting, Relation-R1 achieved dramatic improvements over prior work, reaching 31.67% (Verb) and 3.41% (Grnd-all) compared to CoFormer’s [36] 0.63% and 0.01%, respectively. This advantage is maintained in the 4-shot setting, where Relation-R1 reached 38.02% (Verb) and 5.92% (Grnd-all), significantly surpassing previous methods. These results clearly demonstrate Relation-R1’s superior ability to generalize and perform robust N-ary relation detection, setting a new SOTA for few-shot GSR.
Joint Binary and N-ary Relation Detection.
Table 5: Performance comparison (%) across various cognitive CoT strategies trained on joint binary relation detection and $N$-ary relation detection, separately. Blue indicates metrics without correct verb constraints.
  Binary Relation $N$-ary Relation
Method CoT Recall mRecall Verb Value Value-all Grnd Grnd-all
SFT - 14.89 12.66 56.23 45.81 (66.00) 30.31 (34.06) 38.59 (54.11) 28.77 (36.18)
SFT + RL Template-based 22.33 20.07 57.26 46.66 (66.14) 30.92 (34.75) 40.21 (55.50) 30.18 (37.64)
SFT + RL MLLM-generated 19.59 18.30 53.00 42.27 (65.11) 26.67 (34.65) 35.69 (55.36) 25.89 (37.61)
SFT + RL Progressive 27.19 26.12 73.25 66.18 (82.54) 52.90 (57.83) 55.20 (67.13) 42.36 (48.81)
Beyond training Relation-R1 solely on binary or N-ary relation detection tasks (as discussed in §3), we also investigated the joint training of both tasks within the same Relation-R1 model. The results of this joint training are presented in Table 5333The Grnd-all metric excludes ground-truth bounding boxes represented as [-1, -1, -1, -1] for calculation.. From the analysis of these results, three primary conclusions can be drawn: 1) When utilizing both the SFT and RL strategies, our Relation-R1 consistently achieves higher performance across nearly all metrics for both binary and N-ary relation detection tasks. Similar to single-task training, the MLLM-generated CoT strategy underperforms in the N-ary task. This is likely due to the inherent generality of the CoT generated by Qwen [39], which may not precisely align with the ground-truth annotations, consequently leading to lower Verb accuracy. 2) Our proposed progressive CoT-guided paradigm consistently achieves the highest performance across all evaluated metrics for both tasks in the joint learning setting. 3) The joint progressive CoT-guided training across both binary and N-ary tasks fosters a synergistic mutual promotion effect. This is evidenced by the notable gains observed: for instance, a Recall of 27.19% in Table 5 vs. 22.57% in Table 3 (for binary relation detection), and a Verb accuracy of 73.25% in Table 5 vs. 71.04% in Table 3 (for N-ary relation detection).
5 Qualitative Results
Figure 7: Performance(%) comparison between ASMv2 and Relation-R1 over all predicates on test set of PSG [17] dataset. The orange area denotes the predicate distribution of the training set.
Figure 8: Qualitative results of binary relation detection.
Figure 9: Qualitative results of N-ary relation detection.
Figure 10: Qualitative results of synonymous in N-ary relation detection.
Comparison over All Predicates. We visualized predicate-level comparison against ASMv2 [8] and the predicate distribution in Figure 7. As seen, Relation-R1 outperforms ASMv2 by a clear margin on both prevalent and rare predicates. These results clearly validated the effectiveness and efficiency of our approach for the SGG task in the open-ended setting.
@@ 5 Qualitative Results
Binary and N-ary Relation Detection Results. We visualized some examples of binary and N-ary relation detection in Figure 8 and Figure 9, respectively. As seen, for binary relation detection (c.f. Figure 8), the visualized thinking process reveals how Relation-R1 effectively decomposes the complex task into constituent cognitive steps: object classification, localization, and relation inference. For instance, Relation-R1 can first focus on the main objects (e.g., “cat” and “train”) and then infer their positions and relationships. Similarly, for N-ary relation detection (c.f. Figure 9), Relation-R1 showcases its understanding by first identifying the core action (crowning) and then methodically grounding each associated semantic role – Agent (woman), Object (tiara), and Recipient (woman) – to its corresponding entity and bounding box. This structured analysis underscores Relation-R1’s ability to parse complex events involving multiple participants and their distinct roles within the scene.
Synonymous Visualization. Figure 10 displays synonymous results observed in N-ary relation detection. As evinced by these results, our progressive CoT-guided training paradigm empowers the model to generate diverse synonymous expressions, encompassing entities, semantic roles, and activities. For instance, in Figure 10(top), Relation-R1 demonstrates the ability to output synonymous agents such as “person” and “man”. Furthermore, the model can generate semantically similar and plausible activities, exemplified by “barbecues” and “grills” (c.f. Figure 10(bottom)). Even within the same activity “rests”, it can differentiate and utilize varying semantic role names like “item” and “bodypart” (c.f. Figure 10(row 2)). These synonymous outputs indicate a more flexible and robust semantic understanding beyond rigid lexical matching, thereby significantly enhancing the model’s generalization capabilities to diverse linguistic formulations and real-world variations.
@@ 6 Conclusion
This paper reveals critical limitations in MLLM visual relation understanding, particularly for N-ary relationships and language-induced bias. We propose Relation-R1, a unified framework with a two-stage approach: establishing foundational reasoning and structured outputs via cognitive CoT-guided SFT, followed by RL with GRPO for refinement, promoting visual-semantic grounding and robust generalization. Furthermore, progressively applied cognitive CoT guidance, transitioning from template-based to MLLM-generated reasoning, further enhances generalization, especially for synonymous N-ary relations. Extensive experiments on PSG and SWiG validate Relation-R1’s effectiveness, demonstrating SOTA performance in both binary and N-ary relation detection. Integrating progressive CoT-guided SFT with R1-inspired RL refinement, Relation-R1 advances MLLMs towards more human-like relational reasoning.
Appendix A Implementation Details
We utilized Qwen2.5-VL-3B [39] as our base model. Relation-R1 was trained on an 8xA100 80GB GPU server. During the SFT phase, we employed a total batch size of 32, with 4 samples processed per training step. For the GRPO component, we used a total batch size of 48. The learning rates were set to 2e-5 for SFT and 1e-6 for GRPO, respectively. Qwen2.5-VL-72B [39] served as the CoT generator for producing MLLM-generated CoTs. Progressive training began with supervised fine-tuning on binary and N-ary relation detection tasks for 2 epochs using template-based CoT, followed by fine-tuning on 4k MLLM-generated CoT samples. After that, GRPO RL training was conducted: 2.4k steps on the N-ary relation detection task, 3.6k steps on the binary relation detection task, and finally 2k steps on both tasks.
Appendix B Cognitive CoT Generation Prompt
Question: Scene graph caption is a textual description of an image that incorporates references for objects and their relationships: 1. Use <ref>object class</ref><box>[[x1,y1,x2,y2],…,[x1,y1,x2,y2]]</box> to denote all subject and object regions of a relation triplet in the caption. 2. Use <pred>predicate class</pred><box>subject coordinates list</box><box>object coordinates list</box> to denote all relation triplet regions. 3. Each sentence must have at least one relation triplet region. Given the image, a reasonable scene graph caption is: {ground-truth caption} To make you generate a reasonable scene graph caption, please give the step-by-step reasoning process in <think> </think> tags.
Figure S1: Scene graph caption CoT generation prompt.
Question: Grounded situation frame is a multimodal description combining textual semantics with visual localization information, where key situational elements (agent, action, object, place) are associated with both semantic role labels and corresponding bounding box coordinates in the image. It follows the pattern: <role>entity class</role><box>[x1,y1,x2,y2]</box>. Given the image, a reasonable grounded situation frame is: {ground-truth frame} To make you generate a reasonable grounded situation frame, please give the step-by-step reasoning process in <think> </think> tags.
Figure S2: Grounded situation frame CoT generation prompt.
As mentioned in §3.2, we employ two distinct prompts (Figure S1 and Figure S2) to guide cognitive CoT generation for binary and N-ary relation detection, respectively. Each prompt strategically comprises: 1) a concise task definition outlining the detection objective; 2) an illustrative in-context example demonstrating the desired reasoning steps and output format; and 3) a specific instruction prompting the model to generate CoT of the current image.
Appendix C Scene Graph Format
Question: Generate a scene graph for the provided image as a structured list containing [subject entity, [x1, y1, x2, y2] bounding box coordinates of subject, object entity, [x1, y1, x2, y2] bounding box coordinates of object, relationships between subject and object]. First output the thinking process in <think> </think> tags and then output the final answer in <answer> </answer> tags.
Figure S3: Scene graph format prompt used for training.
The scene graph format prompt employed for GRPO training is displayed in Figure S3. Furthermore, to enable the model’s generation of think process, we utilize templated-CoT (c.f. §3.2) for supervised fine-tuning.
Appendix D Discussions
Limitations. While Relation-R1 demonstrates state-of-the-art performance in visual relation understanding, several limitations warrant discussion and offer avenues for future research. The first one is computational cost. Like other GRPO-based frameworks, the proposed framework introduces a multi-stage training process that requires substantial time to collect a training batch. The second one is the interpretability of GRPO refinements. While CoT aims to make the initial reasoning process more transparent, the subsequent policy optimization by GRPO, while guided by multi-rewards, might still operate as a “black box” to some extent.
Broader Impacts. The development of Relation-R1 promises significant positive impacts, such as more reliable AI systems with reduced hallucinations for applications in robotics and content analysis, enhanced accessibility tools for the visually impaired through richer scene descriptions, and advanced information retrieval based on complex multi-entity interactions. However, ethical considerations include the potential misuse for generating sophisticated misinformation, the risk of amplifying societal biases present in training data, and heightened surveillance capabilities if deployed without oversight.